### v2

# Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation

Abstract
Synthesizing high-quality instruction data from unsupervised text is a promising paradigm for training large language models (LLMs), yet automated methods for this task still exhibit significant limitations in the diversity and difficulty of synthesized instructions. To address these challenges, we propose Self-Foveate, an LLM-driven method for instruction synthesis. Inspired by hierarchical human visual perception, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology that guides the extraction of textual information at three complementary granularities, from fine-grained details through cross-region connections to holistic patterns, thereby enhancing both the diversity and difficulty of synthesized instructions. Furthermore, a re-synthesis module is incorporated to improve the fidelity of instructions to source text and their overall quality. Comprehensive experiments across multiple unsupervised corpora and diverse model architectures demonstrate that Self-Foveate consistently outperforms existing methods. We publicly release our code at https://github.com/Mubuky/Self-Foveate
Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation
Mingzhe Li, Xin Lu, Yanyan Zhaothanks: Email corresponding. Research Center for Social Computing and Interactive Robotics Harbin Institute of Technology, China {mzli, xlu, yyzhao}@ir.hit.edu.cn
Figure 1: Illustration of (a) Self-Foveate in contrast with (b) Baseline Self-QA. For Self-Foveate, the multi-level foveation enables the LLM to extract details (highlighted in distinct colors) of the text, subsequently synthesizing instructions with diversity and difficulty via distinct synthesis paradigms. In comparison, Self-QA employs single-step generation that produces simple and monotonous instruction candidates.
Related Data w/o Human Correctness Diversity Difficulty Work Utilization Annotation Guarantee Augment Augment Self-Instruct (Wang et al., 2023a) Seed QA examples $\checkmark$ Self-Align (Sun et al., 2023) Seed QA examples $\checkmark$ $\checkmark$ Self-Chat (Xu et al., 2023) Dialogue $\checkmark$ Self-QA (Zhang and Yang, 2023) Unsupervised Knowledge $\checkmark$ $\checkmark$ $\checkmark$ LongForm (Köksal et al., 2024) Web dataset $\checkmark$ $\checkmark$ $\checkmark$ Humpback (Li et al., 2024b) Web dataset $\checkmark$ $\checkmark$ ISARA (Guo et al., 2024) Seed QA examples $\checkmark$ $\checkmark$ $\checkmark$ Wiki2023 (Jiang et al., 2024) Unsupervised Text $\checkmark$ $\checkmark$ Self-Foveate (Ours) Unsupervised Text $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: A comparative analysis of various task generation methodologies or frameworks. The gray checkmark symbol denotes that the work may partially accomplish specific objectives (though not comprehensively). The red cross marker indicates either the work’s failure to achieve the stated objective or absence of explicit documentation regarding this goal.
1 Introduction
Large language models (LLMs), such as GPT-4o (OpenAI et al., 2024), Claude 3.5 Sonnet (Anthropic, 2024), and Llama 3.1 (Grattafiori et al., 2024), have demonstrated exceptional instruction-following capabilities (Zhou et al., 2023b) and increasingly strong problem-solving abilities (Cobbe et al., 2021; Dua et al., 2019). A critical component in training such models involves fine-tuning with extensive supervised question-answering instruction data (Wang et al., 2023b). However, substantial challenges persist in constructing large-scale, high-quality supervised fine-tuning (SFT) data for instruction tuning (Zhang et al., 2024). The prohibitive costs of human annotation (Wang et al., 2023a; Sun et al., 2023; Xu et al., 2023), coupled with difficulties in ensuring data diversity and quality control (Ge et al., 2024), continue to impede technological advancements.
Given the strong instruction-following and generative capabilities of modern LLMs, researchers are actively exploring methodologies to leverage these models for synthetic data synthesis (Wang et al., 2023a; Zhang and Yang, 2023; Nayak et al., 2024b; Wu et al., 2024). The primary objectives are to produce high-quality datasets with enhanced diversity and difficulty, thereby improving the performance of fine-tuned models on downstream tasks.
A promising recent paradigm in instruction data synthesis is unsupervised text-based instruction synthesis. The advantages of this paradigm become particularly pronounced when handling massive unsupervised text corpora, a ubiquitous resource containing rich world knowledge and linguistic patterns. By leveraging LLMs’ intrinsic capabilities in contextual understanding and logical reasoning, this paradigm eliminates the need for manual annotation while grounding instruction synthesis in the given unsupervised textual materials.
Recent automated methods such as Self-QA (Zhang and Yang, 2023) have significantly optimized the data generation pipeline while reducing human labor costs. However, as shown in Table 1, critical challenges remain unresolved: (1) Diversity of Instructions: While existing frameworks continuously refine data generation strategies and enhance post-synthesis filtering for instruction data, limitations persist in the diversity of synthesized instructions, which tend to exhibit repetitive patterns in both structure and topic coverage, potentially constraining the generalization of fine-tuned models. (2) Difficulty of Instructions: Current synthesis methodologies generally lack emphasis on instruction complexity and depth. For instance, Self-QA (Zhang and Yang, 2023) directly acquires instructions through single-step generation without guaranteed difficulty levels (as illustrated in Figure 1). Synthesized instructions often exhibit simplistic structures and fail to capture complex inter-entity relationships, limiting models’ ability to produce nuanced responses. These limitations suggest that existing methods fail to fully exploit the rich, multi-granularity information inherent in unsupervised text, leaving substantial potential untapped.
To address these challenges, we observe that textual data inherently contains abundant detailed information encompassing entities (e.g., QuantumX processor), attributes, relations, and writing techniques. For instance, the statement “In the realm of technology, speed is not just a luxury” implicitly frames speed as a fundamental necessity through subtle analogy (as multi-color annotated in Figure 1). This rich, multi-granularity information remains underutilized in existing methods.
This multi-level nature of textual information finds a compelling parallel in human visual perception. In the human visual system, the fovea centralis, a small region of the retina with the highest density of photoreceptor cells, enables sharp central vision for resolving fine details at the point of fixation (Wandell, 1995). Beyond this focal point, rapid saccadic eye movements shift the fovea across different regions of interest, and the brain integrates information from these dispersed fixation points to comprehend spatial relationships (Findlay and Gilchrist, 2003). Meanwhile, peripheral vision captures the global structure and holistic patterns of the scene at a coarser resolution. This multi-level visual processing, spanning fine-grained foveal detail, cross-region saccadic integration, and holistic peripheral awareness, inspires our approach to multi-granularity text comprehension.
Motivated by these textual observations and the analogous multi-level processing in human foveation, this paper proposes Self-Foveate, a comprehensive LLM-based methodology designed to automatically synthesize instructions from unsupervised text. As illustrated in Figure 1, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology in which micro-foveate mirrors foveal fixation on fine-grained details, scatter-foveate emulates saccadic integration across dispersed information, and macro-foveate parallels peripheral perception of holistic patterns, thereby systematically extracting detailed information from raw text and subsequently synthesizing instructions with enhanced diversity and difficulty through three synthesis paradigms. Furthermore, Self-Foveate integrates a re-synthesis module to improve both the fidelity of instructions to source text and their overall quality.
To summarize, the key contributions of this paper are as follows:
  • $\triangleright$
    We systematically analyze unsupervised text-based instruction synthesis, identifying critical limitations of existing methods in both instruction diversity and difficulty.
  • $\triangleright$
    We propose Self-Foveate, a multi-level foveation methodology inspired by human visual perception that synthesizes diverse and challenging instructions from unsupervised text via three complementary strategies: micro-, scatter-, and macro-foveation.
  • $\triangleright$
    We conduct extensive experiments covering diversity and difficulty analysis, downstream task evaluation, and data scaling analysis. Results demonstrate that Self-Foveate consistently outperforms existing methods in unsupervised text-based instruction synthesis.
Figure 2: The Self-Foveate workflow is designed for instruction synthesis based on unsupervised text. Self-Foveate takes unsupervised text as input, extracts foveate elements, foveate groups, and foveate segments, then synthesizes instruction tuning data through these extracted details.
2 Related Work
Instruction Data Synthesis
Multitask instruction fine-tuning (Wei et al., 2022) substantially improves language models’ ability to follow instructions and generalize to unseen tasks (Sanh et al., 2022; Mishra et al., 2022; Chung et al., 2022; Longpre et al., 2023; Zhou et al., 2023a; Li et al., 2024b). As manually curating large-scale instruction data remains costly (Liu et al., 2024) and the constraints that data availability imposes on model capability continue to grow (Villalobos et al., 2024; Long et al., 2024), recent work leverages LLMs to synthesize instruction–response pairs automatically. Self-Instruct (Wang et al., 2023a) bootstraps new tasks from a small set of seed demonstrations; Unnatural Instructions (Honovich et al., 2023) and Alpaca (Taori et al., 2023) scale this idea with minimal human intervention; and Self-QA (Zhang and Yang, 2023) further removes the need for seed tasks by grounding synthesis in unsupervised text. Other efforts refine this pipeline through GPT-4 distillation (Peng et al., 2023), self-rewarding mechanisms (Yuan et al., 2024), and generalized synthesis frameworks (Li et al., 2024a). Our work follows the unsupervised text-based synthesis paradigm, focusing specifically on the quality dimensions of diversity and difficulty that prior pipelines leave underaddressed.
Diversity and Difficulty of Synthesized Instructions
Despite the rapid progress in instruction data synthesis, two persistent challenges remain underexplored. First, synthesized instructions often lack variety in both form and subject matter, narrowing the diversity of the resulting datasets (Chung et al., 2023; Ge et al., 2024). Second, most synthesis pipelines lack explicit mechanisms to control instruction difficulty; for example, Self-QA (Zhang and Yang, 2023) produces instructions in a single pass, offering no explicit control over their complexity. While earlier approaches based on smaller pretrained language models (Schick and Schütze, 2021) were constrained by limited model capacity, even modern LLM-based pipelines have not fully addressed these quality dimensions. A common limitation is that existing methods treat source text as a monolithic input rather than systematically exploiting its multi-granularity structure, leaving substantial room for more varied and challenging instructions. To close this gap, Self-Foveate introduces a multi-level foveation methodology that guides the extraction of textual information at three complementary granularities and yields instructions of greater variety and complexity, all without human annotation.
3 Self-Foveate
We propose Self-Foveate, denoted as $\mathcal{F}$, a multi-level method for synthesizing instructions from unsupervised text without human-annotated samples. Self-Foveate consists of three levels and a re-synthesis module. Formally, given an unsupervised text set $\mathcal{D}$, Self-Foveate transforms each document $d_{i}\in\mathcal{D}$ through a set of synthesis paths $\{\mathcal{F}_{j}\}$. The final synthesized dataset, $\mathcal{D}_{\text{gen}}$, is the union of subsets produced by applying every synthesis path to every document:
$\mathcal{D}_{\text{gen}}=\mathcal{F}(\mathcal{D})=\bigcup_{d_{i}\in\mathcal{D}}\bigcup_{\mathcal{F}_{j}\in\mathcal{F}}\mathcal{F}_{j}(d_{i})$ (1)
The synthesis process is designed to maximize both the diversity and difficulty of the resulting instructions.
As shown in Figure 2, Self-Foveate incorporates three levels and one module:
  • $\triangleright$
    Micro-foveate Level: Extracts fine-grained foveate elements (entities and attributes) from unsupervised text, then uses each element as a seed answer to synthesize instructions via reverse synthesis (§3.1).
  • $\triangleright$
    Scatter-foveate Level: Broadly extracts foveate elements and combines them into foveate groups, then applies direct synthesis to produce instructions requiring cross-entity reasoning (§3.2).
  • $\triangleright$
    Macro-foveate Level: Identifies text segments employing rhetorical or figurative devices as foveate segments, then converts them into instructional form via transcription synthesis (§3.3).
  • $\triangleright$
    Re-synthesis Module: Filters the synthesized instruction set for unanswerable instructions and iteratively re-synthesizes them using successfully synthesized examples as references, improving fidelity to the source text (§3.4).
3.1 Micro-foveate Level
Unsupervised text typically contains a hierarchy of information: primary entities, secondary entities, and their respective attributes. Without explicit guidance, a teacher LLM synthesizing instructions from such text tends to focus on salient surface-level content, overlooking secondary entities and fine-grained attributes. The resulting instruction sets therefore miss critical information, limiting downstream model performance. The micro-foveate level addresses this gap through a fine-grained extraction mechanism paired with reverse synthesis, ensuring that individually important pieces of information are preserved in the synthesized instruction sets. Covering a broader range of entities and attributes enhances diversity, while probing specific entity attributes increases difficulty.
Micro-Foveate Mechanism
This mechanism aims to guide the teacher LLM to synthesize instructions from a fine-grained perspective of the unsupervised text. We introduce the concept of “foveate elements”, which broadly encompass all entities and their attributes within a given text. We guide the teacher LLM to extract more foveate elements than needed and then retain those whose embeddings have the highest cosine similarity to the full text embedding, filtering out extractions that are tangential to the source content.
Reverse Synthesis
Based on the selected foveate elements, we employ the reverse synthesis method to produce instructions. Specifically, we treat each foveate element as a potential answer to an instruction and guide the teacher LLM to synthesize a corresponding instruction from the unsupervised text, then derive the final answer. Even though the foveate elements can already serve as answers or parts of answers, to enhance the fluency, completeness, and accuracy of the answers, we choose to derive the answers anew in the reverse synthesis step rather than directly using the foveate elements as answers.
3.2 Scatter-foveate Level
Beyond individual entities, unsupervised text often encodes implicit relationships (causal links, comparisons, or temporal dependencies) between entities or attributes that are scattered across different passages. A teacher LLM synthesizing instructions without targeted guidance is unlikely to surface these cross-entity connections, producing instruction sets that remain shallow. The scatter-foveate level tackles this by combining dispersed key information through a grouping mechanism and applying direct synthesis, so that the resulting instructions require reasoning about inter-entity relationships. This design broadens diversity by surfacing less obvious implicit information and raises difficulty by demanding cross-entity reasoning.
Scatter-Foveate Mechanism
This mechanism encourages the teacher LLM to reason about connections across different parts of the unsupervised text. We extract foveate elements more broadly from the text and randomly combine them into foveate groups, with group sizes sampled from an empirical distribution.
Direct Synthesis
Based on the formed foveate groups, we synthesize instructions using the direct synthesis method, which treats each element in the foveate group as an indispensable component of the target instruction, guiding the teacher LLM to uncover implicit semantic connections between different entities or attributes and to embed such cross-entity reasoning in the synthesized instructions. After the instructions are synthesized, the corresponding answers are derived from the text.
3.3 Macro-foveate Level
At the broadest granularity, unsupervised text frequently employs figurative language and rhetorical devices (metaphor, hyperbole, contrastive foil, rhetorical questioning, and citation) that convey meaning beyond their literal content. A teacher LLM may overlook these devices without explicit prompting, missing the deeper communicative intent they carry. The macro-foveate level captures this holistic layer by identifying and extracting text segments that employ such writing techniques, then converting them into instructional form through transcription synthesis, ensuring that literary and rhetorical nuances are reflected in the synthesized instruction sets.
Macro-Foveate Mechanism
This mechanism directs the teacher LLM to comprehend unsupervised text from a holistic perspective. We highlight and extract the text segments employing such writing techniques, which are designated as foveate segments.
Transcription Synthesis
Based on the identified foveate segments, we employ transcription synthesis to convert each one into an instructional format. This process transforms declarative foveate segments into interrogative or imperative forms. Subsequently, corresponding answers are synthesized according to the content of the unsupervised text.
3.4 Re-synthesis Module
Due to limitations of the teacher LLM, not all instructions in a directly synthesized instruction set may be answerable from the source text. To improve fidelity, we employ a re-synthesis module that iteratively re-synthesizes these unanswerable instructions over multiple rounds, replacing them with instructions that can be fully answered based on the same text.
Single-Sample Reference Synthesis
During re-synthesis, we process one failed instruction at a time and randomly sample a different subset of successfully synthesized instructions as reference examples in each iteration, improving the success rate by exposing the teacher LLM to varied reference contexts.
Hyperparameter Configuration
The outputs of LLMs are significantly influenced by hyperparameters such as temperature, top-p, and frequency penalty. To improve re-synthesis success rates, we tune these hyperparameters to define a high-creativity configuration, enabling the teacher LLM to synthesize more varied instructions based on reference samples.
Dataset Source # Test Examples Question Context SQuAD Crowdsourced Wikipedia 11639 HotpotQA Crowdsourced Wikipedia 2500 FilmWiki LLM Wikipedia 7398
Table 2: Statistics for the evaluation datasets from our experiments.
Datasets Methods Diversity Metrics SelfBLEU Div. Embedding Div. SQuAD Self-QA 0.593 0.838 Bonito 0.494 0.838 Wiki2023 0.550 0.842 Self-Foveate 0.665 0.851 Test Questions 0.695 0.840 HotpotQA Self-QA 0.463 0.823 Bonito 0.371 0.769 Wiki2023 0.554 0.822 Self-Foveate 0.607 0.835 Test Questions 0.634 0.786 FilmWiki Self-QA 0.406 0.687 Bonito 0.197 0.677 Wiki2023 0.341 0.664 Self-Foveate 0.563 0.706 Test Questions 0.316 0.618
Table 3: Comparison of diversity metrics across different methods and datasets. The table presents SelfBLEU Diversity (SelfBLEU Div.) and Embedding Diversity (Embedding Div.) scores for various methods on the datasets. The diversity of the test questions from each dataset is also provided as a reference.
Model Settings GPT-4o mini DeepSeek-V3 SQuAD HotpotQA FilmWiki SQuAD HotpotQA FilmWiki Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Rec. Acc. Llama-3.1-8B None* 0.309 0.202 0.244 0.160 0.212 0.082 0.309 0.202 0.244 0.160 0.212 0.082 Self-QA 0.367 0.384 0.372 0.358 0.328 0.201 0.389 0.412 0.399 0.378 0.370 0.239 Wiki2023 0.327 0.361 0.338 0.322 0.333 0.235 0.342 0.370 0.340 0.328 0.349 0.244 Bonito* 0.386 0.405 0.360 0.372 0.219 0.153 0.386 0.405 0.360 0.372 0.219 0.153 Self-Foveate 0.484 0.490 0.507 0.486 0.512 0.367 0.481 0.491 0.525 0.501 0.548 0.397 Qwen2.5-7B None* 0.251 0.300 0.266 0.234 0.139 0.032 0.251 0.300 0.266 0.234 0.139 0.032 Self-QA 0.249 0.232 0.276 0.246 0.206 0.082 0.119 0.125 0.102 0.106 0.111 0.056 Wiki2023 0.215 0.221 0.135 0.112 0.192 0.093 0.170 0.083 0.197 0.203 0.202 0.136 Bonito* 0.143 0.109 0.212 0.199 0.168 0.098 0.143 0.109 0.212 0.199 0.168 0.098 Self-Foveate 0.408 0.414 0.372 0.329 0.283 0.140 0.388 0.389 0.342 0.331 0.261 0.140 Gemma-2-9B None* 0.224 0.121 0.175 0.078 0.211 0.099 0.224 0.121 0.175 0.078 0.221 0.099 Self-QA 0.383 0.409 0.408 0.389 0.429 0.315 0.402 0.435 0.424 0.408 0.509 0.386 Wiki2023 0.336 0.378 0.361 0.352 0.478 0.384 0.364 0.399 0.373 0.365 0.494 0.401 Bonito* 0.411 0.457 0.366 0.373 0.255 0.196 0.411 0.457 0.366 0.373 0.255 0.196 Self-Foveate 0.507 0.525 0.537 0.520 0.672 0.528 0.499 0.514 0.552 0.525 0.697 0.581
Table 4: Recall (Rec.) and LLM Accuracy (Acc.) on downstream tasks: Self-Foveate vs. baselines. Results include models fine-tuned with instructions synthesized by GPT-4o mini or DeepSeek-V3, as well as reference non-instruction-tuned models (None). * Indicates that the base model was not fine-tuned using instructions synthesized by GPT-4o mini or DeepSeek-V3.
4 Experiment
Our experiments focus on three critical research questions: (Q1) How effective is Self-Foveate in enhancing instruction diversity? (Q2) How effective is Self-Foveate in enhancing instruction difficulty? (Q3) How effective is Self-Foveate in improving the model’s problem-solving capabilities during instruction fine-tuning?
4.1 Experimental Setup
Datasets
We employ three independent datasets, including the training set of the FilmWiki dataset containing 2,385 unsupervised texts with corresponding question-answer pairs. Additionally, we sample two widely used benchmark QA datasets from the MRQA 2019 shared task (Fisch et al., 2019): SQuAD (Rajpurkar et al., 2016) (following Bonito (Nayak et al., 2024a)) and HotpotQA (Yang et al., 2018). To maintain comparable data scale and computational costs to the FilmWiki dataset, we extract 2,500 unsupervised texts with corresponding QA pairs from each dataset’s training split as substitutes for the complete collections. Further implementation details are provided in Table 2.
Baselines
We consider four baselines: zero-shot, Self-QA, Bonito, and Wiki2023. For the zero-shot baseline, we prompt models for evaluation without leveraging any unsupervised texts from the target task (None). The Self-QA baseline employs an unsupervised knowledge-guided method for extracting instruction-question-answer triples (Self-QA(Zhang and Yang, 2023). The Bonito baseline utilizes a 7B-sized specialized model to generate various types of questions from unsupervised text. We configure it to produce questions of the “question answering without choices” type and obtain a sufficient number of samples through multiple sampling iterations (Bonito). The Wiki2023 baseline implements a text-based QA pair extraction methodology (Wiki2023). All baselines use the same unsupervised texts as our method, as specified in Section 4.1.
Instruction Synthesis
As described in section 3, we process unsupervised texts through Self-Foveate using GPT-4o mini and DeepSeek-V3 to synthesize instructional data. We implement two distinct hyperparameter configurations emphasizing high stability and high creativity, respectively, with detailed specifications in Appendix B.
Base Models
We select three state-of-the-art open-source foundation models as our base models prior to instruction tuning: Meta-Llama-3.1-8B (Grattafiori et al., 2024), Qwen2.5-7B (Qwen et al., 2025), and Gemma-2-9B (Team et al., 2024). These decoder-only language models employ next-word prediction objectives and were pretrained on trillions of tokens without any instruction-based fine-tuning.
4.2 Diversity Analysis
To explicitly evaluate the diversity of the instructions in a metric-driven manner, we follow established practices recommended by Zhu et al. (2018); Perez et al. (2022); Tevet and Berant (2021) and employ two metrics: SelfBLEU scores and Sentence-BERT embedding distances (Reimers and Gurevych, 2019a). These metrics capture different facets of diversity. SelfBLEU measures surface-level textual diversity, while embedding distances measure semantic diversity. For SelfBLEU, we compute average scores using $n$-grams for $n\in\{2,3,4,5\}$, following the approach suggested by Zhu et al. (2018). We use the implementation of the SelfBLEU metric by Alihosseini et al. (2019). Further details are available in Appendix G.
Table 3 demonstrates that Self-Foveate-synthesized instructions achieve substantial diversity improvements in both textual and semantic dimensions, attaining or even surpassing the diversity level of crowdsourced test questions through a low-cost automated process.
4.3 Difficulty Analysis
To investigate the difficulty of the synthesized instructions, we design prompts and employ GPT-4o to conduct a rigorous head-to-head comparison between the instructions synthesized by Self-Foveate and those synthesized by baseline methods. Specifically, for each unsupervised text in each dataset, we use two contrasting methods to synthesize an equal number of instructions. We then provide GPT-4o with a set of instructions synthesized by both methods under the same unsupervised text, ensuring that the instruction sets are anonymous and their relative positions are randomized. Subsequently, we record GPT-4o’s judgments and calculate the win rate for Self-Foveate. The results of the head-to-head comparison, as shown in Table 5, demonstrate the consistently higher difficulty of the instructions synthesized by Self-Foveate. Further details can be found in Appendix H.
Figure 3: Impact of instruction set scale on model fine-tuning performance for Self-Foveate and baselines: (a) Recall and (b) LLM Accuracy.
Dataset Baseline Self-Foveate WR. Win Lose SQuAD Self-QA 70.64% 29.36% Wiki2023 80.83% 19.17% Bonito 99.96% 00.04% HotpotQA Self-QA 89.52% 10.48% Wiki2023 91.17% 08.83% Bonito 100.00% 00.00% FilmWiki Self-QA 85.12% 14.88% Wiki2023 95.08% 04.92% Bonito 96.31% 03.69%
Table 5: Head-to-head comparison of instruction difficulty for Self-Foveate against baselines through win rates (WR.) across datasets.
4.4 Problem-solving Capabilities
To evaluate the impact of instruction-tuning datasets on model performance in downstream tasks, we fine-tune open-source models and assess their performance on in-distribution problems. Table 4 illustrates the fine-tuning performance of open-source models using instruction sets synthesized by two teacher LLMs (GPT-4o mini and DeepSeek-V3), with equal quantities of instructions per unsupervised text. As answers in test problems often consist of phrase-level or sentence-fragment extractions from source texts while large language models (LLMs) tend to generate more comprehensive responses, we select both recall and LLM-evaluated accuracy as evaluation metrics. Higher recall indicates that generated answers are more likely to contain correct responses, while higher LLM-evaluated accuracy reflects stronger agreement between model-produced answers and ground truth labels, as judged by an LLM evaluator.
To control potential confounding effects from the number of instructions synthesized per unsupervised text and examine the agreement between these two metrics, we investigate the influence of the scale of instruction sets synthesized by Self-Foveate and baseline methods on model fine-tuning performance using the HotpotQA dataset, as shown in Figure 3. Additional details regarding downstream task evaluations are provided in Appendix I.
Table 4 demonstrates that all three base models achieve the best performance across downstream tasks in both metrics when fine-tuned with instructions synthesized via Self-Foveate, regardless of the teacher LLM used for synthesis. Notably, for Qwen2.5-7B, instruction data synthesized by baseline methods can lead to performance degradation compared to the base model without fine-tuning, further demonstrating the superiority of Self-Foveate in instruction synthesis. Our analysis in Figure 3 reveals that as the number of instructions synthesized by Self-Foveate per unsupervised text increases, downstream performance improves significantly, with an expanding gap compared to baseline methods. Figure 3 also confirms the consistency between recall and LLM-evaluated accuracy across varying instruction set scales.
4.5 Ablation Studies
To validate the effectiveness and necessity of each component in Self-Foveate, we conduct comprehensive ablation studies focusing on two key aspects: (1) the contribution of each core component in the multi-level foveation framework, and (2) the impact of the answer regeneration mechanism in reverse synthesis.
4.5.1 Component-wise Analysis
We evaluate the necessity of each core component by systematically removing individual components from the complete Self-Foveate framework. Table 6 presents the ablation results on the SQuAD dataset using Llama-3.1-8B as the base model.
Setting Recall LLM Acc. w/o Micro-Foveate 0.283 0.277 w/o Scatter-Foveate 0.274 0.260 w/o Macro-Foveate 0.344 0.339 Self-Foveate (Full) 0.484 0.490
Table 6: Ablation study results showing the contribution of each core component in Self-Foveate on the SQuAD dataset with Llama-3.1-8B.
The removal of any single component leads to clear performance degradation, with scatter-foveate removal causing the largest drop (Recall from 0.484 to 0.274), confirming that all three components are necessary and complementary.
4.5.2 Answer Regeneration Analysis
In our reverse synthesis paradigm, although foveate elements (e.g., “QuantumX processor”) can directly serve as answers, we regenerate answers to enhance fluency, completeness, and semantic coherence. To validate this design choice, we compare the quality of answers with and without the regeneration mechanism. Table 7 presents the comparative analysis on the SQuAD dataset, evaluating both fluency and completeness of synthesized answers using GPT-4o.
Setting High (%) Medium (%) Low (%) Fluency w/o Regeneration 64.9 19.5 15.6 Reverse Synthesis 93.0 6.8 0.2 Completeness w/o Regeneration 36.7 27.5 35.7 Reverse Synthesis 76.6 19.7 3.7
Table 7: Comparison of answer quality with and without regeneration mechanism in reverse synthesis on the SQuAD dataset. Fluency and completeness are evaluated on a three-point scale (High/Medium/Low).
The results show that answer regeneration substantially improves both fluency and completeness. This validates our choice to maintain consistency across all synthesis paradigms by synthesizing contextually appropriate and semantically coherent answers from the source text, rather than directly using extracted foveate elements.
5 Conclusion
In this paper, we presented Self-Foveate, an LLM-driven methodology for synthesizing diverse and challenging instructions from unsupervised text. Drawing inspiration from hierarchical human visual perception, Self-Foveate introduces a “Micro-Scatter-Macro” multi-level foveation methodology: micro-foveate captures fine-grained entity details, scatter-foveate integrates dispersed cross-entity relationships, and macro-foveate extracts holistic rhetorical patterns. A re-synthesis module further enhances fidelity to source text and overall instruction quality. Comprehensive experiments across multiple unsupervised corpora and diverse model architectures demonstrate that Self-Foveate consistently outperforms existing methods in instruction diversity, difficulty, and downstream task performance. Our work establishes multi-level foveation as an effective paradigm for high-quality instruction synthesis from unsupervised text.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) via grant 62441614 and 62176078. We thank the anonymous reviewers for their insightful comments and suggestions.
Limitations
Although Self-Foveate provides significant improvements in synthesizing diverse and difficulty-oriented fine-tuning instruction data based on unsupervised text, several limitations must be acknowledged. First, the computational cost of processing large-scale unsupervised text using closed-source state-of-the-art large language models (LLMs) remains substantial. Ensuring data quality requires multi-step reasoning and iterative processing on unsupervised text, which further escalates computational demands. Second, although Self-Foveate aims to fully synthesize data from unsupervised text and incorporates verification mechanisms, the inherent tendency of LLMs to hallucinate or generate erroneous information persists as a challenge. While precise prompting strategies and higher-quality unsupervised text can mitigate these inaccuracies, they cannot be entirely eliminated.
Ethical Consideration
We state that any research or application arising from this study is strictly authorized solely for research purposes. In our work, any unsupervised text datasets used are from public sources and do not contain any private information. In this paper, we have fully presented the prompts used by Self-Foveate in the Appendix. All synthesized instructions rely on the provided unsupervised text and are inspected by relevant modules. Therefore, our method strives to minimize potential safety and ethical risks as much as possible. However, during the process of synthesizing fine-tuning instruction data, maliciously provided unsupervised text can lead the model to produce harmful or inappropriate outputs, which is a shared problem. Additionally, potential unfairness and discrimination present in the unsupervised text might be amplified by LLMs during the instruction synthesis process. Ensuring the quality of synthesized fine-tuning instruction data in a safe and highly controllable manner is crucial. The application of these techniques should be guided by ethical considerations, with safeguards in place to prevent misuse and reduce the likelihood of producing harmful outcomes.
Appendix A Datasets
A.1 Dataset Details
This section provides comprehensive details regarding the datasets utilized in this study.
SQuAD(Rajpurkar et al., 2016)
This Dataset serves as a benchmark for extractive question answering, containing over 100,000 human-generated question-answer pairs anchored to Wikipedia passages. In our work, we utilize this dataset from the MRQA 2019 shared task(Fisch et al., 2019). Considering computational costs, we employ 2,500 randomly selected articles or contexts from the training set.
HotpotQA(Yang et al., 2018)
The original purpose of proposing this dataset was to challenge models with multi-hop reasoning across 113k Wikipedia-based QA pairs requiring synthesis of information from multiple documents. In our work, we utilize this dataset from the MRQA 2019 shared task(Fisch et al., 2019). Considering computational costs, we employ 2,500 randomly selected articles or contexts from the training set.
FilmWiki(Saito et al., 2024)
This dataset was initially constructed to investigate a phenomenon called the perplexity curse, with all texts sourced from Wikipedia. Our study employs its film-themed subset containing the complete collection of 2,385 unsupervised texts and their corresponding questions.
A.2 Usage of Datasets
During the synthesis process, we use the articles or contexts from the dataset (excluding the questions) and synthesize instructions through the Self-Foveate. In the evaluation of problem-solving capabilities, we assess the fine-tuned models using the corresponding questions in the dataset.
Appendix B Instruction Synthesis Details
B.1 Teacher Models and Sampling Hyperparameters
Taking into account both performance and API costs, and to eliminate potential biases in experimental results caused by specific teacher models (such as prompt hacking), we employ two models, GPT-4o mini(gpt-4o-mini-2024-07-18)(OpenAI et al., 2024) and DeepSeek-V3(DeepSeek-AI et al., 2025), to synthesize instructions from unsupervised text using the same methodology and sampling hyperparameters. Detailed hyperparameters are provided in Table 8, where, as described in Section 3, the normal mode is utilized for the initial foveate synthesis, and the high-creativity mode is adopted for re-synthesis.
Hyperparameters Values Normal Mode frequency_penalty 0.5 max_completion_tokens None presence_penalty 0 temperature 0.5 top_p 1.0 High-creativity Mode frequency_penalty 0.5 max_completion_tokens None presence_penalty 0 temperature 1.2 top_p 1.0
Table 8: Hyperparameters of Synthetic Instruction Synthesis.
B.2 Adaptive Task Count Determination
To adaptively determine the number of instructions to synthesize for each article, Self-Foveate employs a two-stage process: (1) counting contextually relevant words via semantic similarity thresholding, and (2) normalizing the resulting counts using a Box-Cox transformation. This ensures that articles with richer content receive proportionally more instructions while maintaining a balanced distribution across the dataset.
Context Thresholding.
For each article $d_{i}$, we compute the text-level embedding $\mathbf{e}_{\text{text}}=\text{Embed}(d_{i})$ and the word-level embeddings $\mathbf{e}_{w_{j}}=\text{Embed}(w_{j})$ for each unique word $w_{j}$ in $d_{i}$. A word is considered contextually relevant if its cosine similarity with the full text exceeds the threshold $\tau$:
$c_{i}=\left|\left\{w_{j}:\text{CosSim}(\mathbf{e}_{\text{text}},\mathbf{e}_{w_{j}})\geq\tau\right\}\right|$ (2)
where $\tau=0.16$ (CONTEXT_THRESHOLD) by default.
Box-Cox Normalization.
The raw counts $\{c_{i}\}$ are then normalized using the Box-Cox transformation to produce the final task counts. The procedure is detailed in Algorithm 1.
Algorithm 1 Adaptive Task Count via Box-Cox Normalization
1:Article set $\mathcal{D}=\{d_{1},\ldots,d_{N}\}$
2:   Threshold $\tau=0.16$, target mean $\mu=8.0$
3:   Scaling factor $\alpha=1.0$
4:Task counts $\{n_{i}\}$ for each article $d_{i}$
5:for each article $d_{i}\in\mathcal{D}$ do
6:  $\mathbf{e}_{\text{text}}\leftarrow\text{Embed}(d_{i})$
7:  $W\leftarrow\text{UniqueWords}(d_{i})$
8:  for each word $w_{j}\in W$ do
9:   $\mathbf{e}_{w_{j}}\leftarrow\text{Embed}(w_{j})$
10:   $s_{j}\leftarrow\text{CosSim}(\mathbf{e}_{\text{text}},\mathbf{e}_{w_{j}})$
11:  end for
12:  $c_{i}\leftarrow|\{w_{j}:s_{j}\geq\tau\}|$
13:end for
14:// Ensure positivity for Box-Cox
15:if $\min(\{c_{i}\})\leq 0$ then
16:  $c_{i}\leftarrow c_{i}+|\min(\{c_{i}\})|+1,\quad\forall\,i$
17:end if
[…]
21:return $\{n_{i}\}$
B.3 Scatter Keyword Grouping
The scatter-foveate level synthesizes instructions from keyword groups of varying sizes (1, 2, or 3 keywords per group). Given a target number of scatter-level instructions $n_{\text{scatter}}$ for an article, the number of groups at each size is determined by two configuration ratios, as described in Algorithm 2.
Algorithm 2 Scatter Keyword Grouping
1:Target instruction count $n_{\text{scatter}}$, ratios $r_{a}=2.50$, $r_{b}=2.25$
2:Keyword groups $\mathcal{G}$
3:// Compute group distribution
[…]
7:// Compute total keywords needed
[…]
9:// Extract and rank keywords via LLM
[…]
13:// Construct groups
[…]
17:return $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}\cup\mathcal{G}_{3}$
The ratios $r_{a}$ (SCATTER_RATIO_A =2.50=2.50) and $r_{b}$ (SCATTER_RATIO_B =2.25=2.25) are designed to ensure a balanced distribution of keyword group sizes. Larger groups (3 keywords) produce more complex, multi-aspect instructions, while single-keyword groups yield focused, targeted questions. By including important (core and major) keywords into the shuffled pool, we ensure that critical concepts appear across multiple group sizes, further enhancing instruction diversity.
B.4 Re-synthesis Module
The re-synthesis module handles cases where initial instruction generation fails, either due to invalid format or unanswerable content. The process follows a structured retry mechanism:
Validation Criteria.
A generated instruction-answer pair is considered valid if it satisfies all of the following:
  1. 1.
    The output matches the expected “Question: ... Answer: ...” format (verified via regex parsing).
  2. 2.
    The answer does not contain “I don’t know,” indicating the model could not derive an answer from the background knowledge.
  3. 3.
    The instruction and answer are both non-empty after stripping whitespace.
Re-synthesis Flow.
For each failed instruction, the re-synthesis module applies the retry procedure described in Algorithm 3.
Algorithm 3 Re-synthesis Retry Procedure
1:Failed index set $\mathcal{F}$, successful pairs $\mathcal{S}$, article $d$, max retries $R=5$
2:Re-synthesized instruction set $\mathcal{I}_{\text{new}}$
3:for each failed index $i\in\mathcal{F}$ do
4:  $f_{i}\leftarrow\text{GetFeature}(i)$ $\triangleright$ keyword / group / sentence
5:  for $r=1$ to $R$ do
6:   $\mathcal{E}\leftarrow\text{Shuffle}(\mathcal{S})$ $\triangleright$ Few-shot examples
7:   $(q,a)\leftarrow\text{LLM}(d,\;f_{i},\;\mathcal{E})$
8:   if $\text{IsValid}(q,a)$ then $\triangleright$ Format + answerable + non-empty
9:     $\mathcal{I}_{\text{new}}\leftarrow\mathcal{I}_{\text{new}}\cup\{(q,a)\}$
10:     break
11:   end if
12:  end for
13:end for
14:return $\mathcal{I}_{\text{new}}$
B.5 Complete Hyperparameter Summary
Table 9 provides a complete summary of all framework-level hyperparameters used in the Self-Foveate pipeline.
Parameter Value Description Thresholding & Normalization CONTEXT_THRESHOLD 0.16 Cosine similarity threshold for context relevance DEFAULT_TARGET_MEAN ($\mu$) 8.0 Target mean for Box-Cox normalization DEFAULT_ALPHA ($\alpha$) 1.0 Scaling factor for Box-Cox transformation Scatter Keyword Grouping SCATTER_RATIO_A 2.50 Ratio for single-keyword groups SCATTER_RATIO_B 2.25 Ratio for two-keyword groups Processing Configuration DEFAULT_MAX_RETRIES 5 Maximum re-synthesis attempts DEFAULT_MAX_WORKERS 96 Parallel processing threads (articles) EMBEDDING_MAX_WORKERS 10 Parallel threads for embedding generation
Table 9: Complete hyperparameter configuration of the Self-Foveate framework.
Appendix C Prompt Templates – Micro-foveate Level
This section presents all prompt templates used in the Micro-foveate level of the Self-Foveate pipeline. The micro-level strategy extracts individual keywords from the article and synthesizes instructions based on those keywords.
C.1 Keywords Extraction
Micro Keywords Extraction – System Prompt You are a professional text analysis expert. Your task is to extract keywords from the given article. Based on the following article, please extract {num_micro} semantically different keywords in random order, with: Please ensure: 1. All keywords are based on the article content. 2. Each keyword is unique to maintain diversity. 3. Keywords can be single words or meaningful phrases. 4. Keywords should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. keyword1 2. keyword2
Micro Keywords Extraction – User Prompt Article content: {content}
C.2 Instruction Generation
Micro Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_micro} questions/instructions based on the given background knowledge and answers. Based on the following background knowledge and answers, please generate one question/instruction for each answer. Requirements: 1. Each generated question must be answerable by its corresponding answer 2. Questions/instructions must be based on the background knowledge content 3. Each question/instruction should focus on different aspects or information from the background knowledge 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Micro Instruction Generation – User Prompt The background knowledge is: {content} The answers are: {answers} Please generate {num_micro} questions/instructions based on the background knowledge and answers above.
C.3 Instruction Regeneration
Micro Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target answer. Based on the following background knowledge and answers, please generate one question/instruction for the target answer. Requirements: 1. The generated question/instruction must be answerable by its corresponding answer. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Micro Instruction Regeneration – User Prompt The background knowledge is: {content} The target answer is: {answer}
Appendix D Prompt Templates – Scatter-foveate Level
This section presents all prompt templates used in the Scatter-foveate level of the Self-Foveate pipeline. The scatter-level strategy extracts keyword groups of varying sizes (1, 2, or 3 keywords per group) and synthesizes instructions that incorporate all keywords from each group.
D.1 Keywords Extraction
Scatter Keywords Extraction – System Prompt You are a professional text analysis expert. Your task is to extract keywords from the given article. Based on the following article, please extract {num_keywords} semantically different keywords in random order, with: Please ensure: 1. All keywords are based on the article content. 2. Each keyword is unique to maintain diversity. 3. Keywords can be single words or meaningful phrases. 4. Keywords should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. keyword1 2. keyword2
Scatter Keywords Extraction – User Prompt Article content: {content}
D.2 Important Keywords Selection
Scatter Important Keywords Selection – System Prompt You are a professional text analysis expert. Your task is to identify the most important keywords from the given background knowledge. Please analyze the given keywords and select: 1. {num_core} core keywords that are absolutely critical (primary keywords) 2. {num_major} major keywords that are highly important but secondary (secondary keywords) For selecting the core keyword (select exactly {num_core}): - Must be the most essential concept that captures the central theme - Should be impossible to understand the text without this keyword - Removing this keyword would make the text lose its main focus For selecting major keywords (select exactly {num_major}): - Should be key supporting concepts that elaborate on the core keyword - Must be directly relevant to explaining or contextualizing the main topic - Should cover different important aspects of the content Please output strictly in this format without any other text: Core keyword: 1. [first core keyword] 2. [second core keyword] Major keywords: 1. [first major keyword] 2. [second major keyword]
Scatter Important Keywords Selection – User Prompt Background knowledge: {content} Keywords: {keywords}
D.3 Instruction Generation
Scatter Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_scatter} questions/instructions based on the given background knowledge and keyword groups. Based on the following background knowledge and keyword groups, please generate one question/instruction for each keyword group that incorporates ALL keywords from that group. Requirements: 1. Each generated question/instruction must use ALL keywords from its corresponding group 2. Questions/instructions must be based on the background knowledge content 3. Each question/instruction should focus on different aspects or information from the background knowledge 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Scatter Instruction Generation – User Prompt The background knowledge is: {content} The keyword groups are: {keyword_groups} Please generate {num_scatter} questions/instructions based on the background knowledge and keyword groups above.
D.4 Instruction Regeneration
Scatter Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target keyword group that incorporates ALL keywords from that group. Based on the following background knowledge and keyword groups, please generate one question/instruction for the target keyword group that incorporates ALL keywords from that group. Requirements: 1. The generated question/instruction must use ALL keywords from its corresponding group. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Scatter Instruction Regeneration – User Prompt The background knowledge is: {content} The target keyword group to use: {keyword_group}
Appendix E Prompt Templates – Macro-foveate Level
This section presents all prompt templates used in the Macro-foveate level of the Self-Foveate pipeline. The macro-level strategy extracts complete sentences as contextual features and synthesizes instructions based on those sentences.
E.1 Sentence Extraction
Macro Sentence Extraction – System Prompt You are a professional text analysis expert. Your task is to extract sentences from the given article. Based on the following article, please extract {num_macro} semantically different key sentences in random order, with: Please ensure: 1. All key sentences are based on the article content. 2. Each key sentence is unique to maintain diversity. 3. Key sentences should focus on identifying key details, focusing on transitions, quotations, comparisons, and rhetorical devices. 4. Key sentences should cover the main entities in the text and their related attributes or related entities. 5. Output strictly follows the format below without any other text: 1. key sentence1 2. key sentence2
Macro Sentence Extraction – User Prompt Article content: {content}
E.2 Instruction Generation
Macro Instruction Generation – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate {num_macro} questions/instructions based on the given background knowledge and key sentences. Based on the following background knowledge and key sentences, please generate one question/instruction for each key sentence. Requirements: 1. Each generated question/instruction must use the corresponding key sentence. 2. Questions/instructions must be based on the background knowledge content. 3. Each question/instruction should focus on different aspects or information from the background knowledge. 4. Output strictly follows JSON format as shown below without any other text: {json_data}
Macro Instruction Generation – User Prompt The background knowledge is: {content} The key sentences are: {sentences} Please generate {num_macro} questions/instructions based on the background knowledge and key sentences above.
E.3 Instruction Regeneration
Macro Instruction Regeneration – System Prompt You are an excellent questions/instructions generation expert. Your task is to generate one question/instruction for the target key sentence. Based on the following background knowledge and key sentences, please generate one question/instruction for the target key sentence. Requirements: 1. The generated question/instruction must use the corresponding key sentence. 2. The question/instruction must be based on the background knowledge content. 3. The question/instruction should be fully answerable using only the background knowledge. 4. Output the generated question/instruction only without any other text.
Macro Instruction Regeneration – User Prompt The background knowledge is: {content} The target key sentence to use: {sentence}
Appendix F Prompt Templates – Task Generation
This section presents all prompt templates used in the Task Generation module of the Self-Foveate pipeline. This module handles reading comprehension answer generation and answer improvement through rereading.
F.1 Reading Comprehension (Answer Generation)
Reading Comprehension (Answer Generation) – System Prompt You are an excellent reading comprehension assistant designed to help users answer the question based on provided background knowledge. Please carefully read the background knowledge and accurately answer the question posed based on its content. - Do not alter key information from the original text. - Avoid using phrases like “based on the above article.” - If the background knowledge lacks sufficient information to answer a question, respond with “I don’t know.” - Ensure your answers are derived from the background knowledge, detailed, and accurate. - Answer the question thoroughly. Please generate the answer following this format strictly: Question: [question] Answer: [detailed and accurate answer]
Reading Comprehension (Answer Generation) – User Prompt The background knowledge is: {unsupervised_knowledge_data} Please answer the following question based on the content of the article above: {instruction}
F.2 Answer Rereading (Answer Improvement)
Answer Rereading (Answer Improvement) – System Prompt You are a smart AI assistant. For a given question-answer pair, improve the answer by correcting errors, bolstering informativeness, aligning with the question, and providing comprehensive detail. - Do not alter key information from the original text. - Avoid using phrases like “based on the above article.” - Ensure your answers are derived from the background knowledge, detailed, and accurate. - Answer the question thoroughly. Please generate the rewritten answer following this format strictly: Question: [question] Answer: [detailed and accurate rewritten answer]
Answer Rereading (Answer Improvement) – User Prompt The background knowledge is: {unsupervised_knowledge_data} The question is: {instruction} The original answer is: {output} Please reread the background knowledge and generate a more detailed and accurate rewritten answer based on the content of the article above.
Appendix G Diversity Experiment
Mathematically, we define the diversity metrics as follows (’D’ denotes ’Diversity’, ’SB’ denotes ’SelfBLEU’, and ’EB’ denotes ’Embedding’):
$\text{D}_{\text{SB}}$ $=1-\dfrac{1}{\lvert\mathcal{X}_{\tau}\rvert}\sum_{x_{i}\in\mathcal{X}_{\tau}}\sum^{5}_{n=2}\text{SelfBLEU}_{\mathcal{X}_{\tau}}(x_{i},n)$ (3)
$\text{D}_{\text{EB}}$ $=1-\dfrac{2}{\lvert\mathcal{X}_{\tau}\rvert(\lvert\mathcal{X}_{\tau}\rvert-1)}\sum_{i<j}\dfrac{\phi(x_{i})\cdot\phi(x_{j})}{\lVert\phi(x_{i})\rVert\lVert\phi(x_{j})\rVert}$ (4)
For computing sentence embeddings, we employ sentence-transformers/paraphrase-multilingual-mpnet-base-v2(Reimers and Gurevych, 2019b), which maps sentences to a 768-dimensional dense vector space. Compared to other sentence embedding transformation models, it offers a balanced mapping dimension and accuracy, enabling the computation of cosine similarity matrices at an acceptable computational cost while obtaining relatively accurate diversity evaluation metrics.
Appendix H Difficulty Head-to-Head Experiment
We prompt GPT-4o (gpt-4o-2024-08-06)(OpenAI et al., 2024) to determine which method-synthesized instruction set is more challenging. The prompts are shown below.
Difficulty Head-to-Head Experiment System Prompt Please compare two sets of instructions/questions for a given content to determine which set is more difficult for a language model to generate coherent and accurate responses. A set of instructions is considered more difficult if it requires: 1. More complex reasoning or multi-step analysis 2. Integration of multiple concepts or contexts 3. Explanation of nuanced relationships or processes 4. Handling of ambiguous or less-documented information 5. Synthesis of information from diverse sources or time periods Please respond with ’1’ if the first set of instructions is more difficult, or ’2’ if the second set of instructions is more difficult. Please respond with the number only, no other text or characters.
Difficulty Head-to-Head Experiment User Prompt Content: {unsupervised text} Instruction Set 1: {instruction set a} Instruction Set 2: {instruction set b}
The hyperparameters set during sampling are listed in Table 10.
Hyperparameters Values frequency_penalty 0.0 max_completion_tokens 5 presence_penalty 0 temperature 0.5 top_p 1.0
Table 10: Hyperparameters of Difficulty Head-to-Head Experiment.
Appendix I Downstream Experiment
I.1 Software and Hardware Details
The implementation leverages the LLaMA-Factory framework (Zheng et al., 2024) with computational optimizations from FlashAttention2 (Dao et al., 2022; Dao, 2024) and Unsloth libraries (Daniel Han and team, 2023). For training, we employ NVIDIA A100, A800, and H800 GPUs based on availability within our computational cluster. The entire work required approximately 110 GPU hours.
I.2 Hyperparameters
The hyperparameters and technical configurations for the instruction tuning process are documented in Table 11.
Hyperparameters Values cutoff_len 2048 learning_rate 0.0001 num_train_epochs 5.0 effective_batch_size 16 lr_scheduler_type cosine max_grad_norm 1.0 warmup_steps 0 optim adamw_torch quantization_bit 4 quantization_method bitsandbytes lora_rank 8 lora_alpha 16 lora_dropout 0
Table 11: Hyperparameters of Instruction Tuning with Q-LoRA Quantization.
I.3 LLM Accuracy Evaluation
LLM Accuracy Evaluation System Prompt You are a fair judge. Your task is to determine if the generated answer correctly answers the question, even if it contains additional explanations. Rules: 1. The generated answer is correct if it contains the key information from the ground truth 2. Additional explanations or context in the generated answer should not make it incorrect 3. Only respond with ’Correct’ or ’Incorrect’
LLM Accuracy Evaluation User Prompt Compare the following answers: Question: {question} Ground Truth Answer: {ground_truth} Generated Answer: {generated} Is the generated answer correct, regardless of any additional explanation? Respond only with ’Correct’ or ’Incorrect’.
Appendix J Answer Regeneration Evaluation
This section provides detailed documentation of the evaluation methodology used in the ablation study for answer regeneration analysis (Section 4.5). To assess the quality of answers generated with and without the regeneration mechanism, we employed GPT-4o (gpt-4o-2024-08-06) to evaluate fluency and completeness on a three-point scale.
J.1 Evaluation Prompt
The following system prompt was used to evaluate question-answer pairs:
Answer Regeneration Evaluation System Prompt Please evaluate the given question and answer pair based on two criteria: 1. Fluency: How well does the answer flow and connect with the question? 1. - High: The answer naturally follows from the question 1. - Medium: The connection is somewhat clear but could be improved 1. - Low: The answer feels disconnected from the question 2. Completeness: How thoroughly does the answer address the question? 2. - High: The answer fully addresses all aspects of the question 2. - Medium: The answer covers most aspects but misses some points 2. - Low: The answer only partially addresses the question Please respond with a JSON object in the following format: { …."fluency": "high|medium|low", …."completeness": "high|medium|low" } Do not include any other text or explanation.
J.2 Evaluation Hyperparameters
The hyperparameters used for the answer regeneration evaluation are provided in Table 12.
Hyperparameters Values frequency_penalty 0.0 max_completion_tokens 50 presence_penalty 0 temperature 0.2 top_p 1.0
Table 12: Hyperparameters for Answer Regeneration Evaluation.
Appendix K Use of AI Assistants
In this study, we utilized AI-powered tools, including ChatGPT, to improve the linguistic quality of the manuscript through spell-checking and minor grammatical corrections. Additionally, the codebase was developed using Cursor to enhance coding efficiency and accuracy.